OpenStudy (anonymous):

The price of an electronic puzzle was reduced by $7. Mr. Murray bought 10 puzzles for his relatives. The total cost before taxes was $130. What was the original price of the puzzle?

OpenStudy (yesyesyes):

20 dollars

OpenStudy (yesyesyes):

130/10 =13 + 7 = 20

OpenStudy (anonymous):

thanks!!

OpenStudy (mathteacher1729):

Hint 1: If he bought 10 of them for $130 what is the cost of just one?